# On nondegenerate $\mathbb{Z}_{2}$-harmonic $1$-forms with shrinking branching sets

### v1

@@ Abstract.
We develop a gluing theorem for non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms on compact manifolds, in which non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms on $\mathbb{R}^{n}$ are glued to the regular zeros of a non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-form. As an immediate consequence, viewing an ordinary harmonic $1$-form as a $\mathbb{Z}_{2}$-harmonic $1$-form without branching set, we prove that for every compact manifold $M^{n},n\geq 3$, if the first Betti number $b^{1}(M)>0$, then $M$ admits a family of non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms, which resolves a folklore conjecture. We will also discuss several possible applications to special holonomy, in particular, to the field of $G_{2}$-geometry.
@@ 1. Introduction
$\alpha=\mathrm{d}\bigg(\mathrm{Re}(A(t)\zeta^{\frac{1}{2}}+B(t)\zeta^{\frac{3}{2}})-\frac{1}{2}\mathrm{Re}(\bar{\mu}\zeta)\mathrm{Re}(A(t)\zeta^{\frac{1}{2}})+O(r^{\frac{5}{2}})\bigg),$
@@
The key feature in Theorem 1.3 is that the Riemannian metric $g$ remains invariant. This is crucial for geometric applications in special holonomy. On the other hand, if this condition is dropped, one can apply a modification of Calabi’s intrinsic method [2] to construct a family of metrics $g^{\prime}_{\epsilon}$ such that the closed $1$-form obtained in the pre-gluing in Section 3.2 is harmonic with respect to those new metrics. And moreover, in this case, the non-degenerate condition for the $\mathbb{Z}_{2}$-harmonic $1$-form on $\mathbb{R}^{n}$ can be dropped. For more related discussion regarding Calabi’s intrinsic method, see Section 5.1.
@@
The author’s main motivation for studying this problem comes from a non-linear version developed by Donaldson in [9], involving the branched maximal surface equation and the adiabatic limit of Kovalev-Lefchetz coassociative fibration. Let $\mathbb{R}^{3,19}=H^{2}(K3,\mathbb{R})$ and let $\Lambda=H^{2}(K3,\mathbb{Z})$ be the integral lattice. Also let $\Sigma$ be a smooth link in a compact $3$-manifold $B$ and let $\Gamma$ be the affine extension
@@
$h=(t,\zeta,w,\mathrm{Re}(B(t)\zeta^{\frac{3}{2}}))+O(|\zeta|^{\frac{5}{2}}),$
near $\Sigma$. Here, $t$ is the coordinate on $\Sigma$ and $\zeta$ is the complex coordinate in the normal direction. We call $h$ avoids (2)(-2)-classes if away from the branching set $\Sigma$ there is no (2)(-2)-class orthogonal to the image of the derivatives of $h$. In [9], Donaldson conjectured that if there is a branched maximal section $h$ that avoids (2)(-2)-classes, then there is a compact $7$-manifold $X$ admitting a differentiable Kovalev-Lefschetz fibration $\pi:X\to B$ (Definition 1, page 16 [9]) with generic fibers diffeomorphic to a $K3$ surface, and there is a one parameter family of torsion-free $G_{2}$-structures $\varphi_{t},t>t_{0}\gg 1$ on $X$ such that the fibers of $\pi$ are coassociative. The singular locus of the fibration coincides with $\Sigma$. Moreover, the normalized hyper-symplectic structures on the $K3$ fibers will converge to a family of hyper-Kähler structures induced by the differential of the branched maximal section $h$ at each $b\in B$. Constructing a solution to the branched maximal surface equation over a compact base space is challenging, particularly under the additional requirement that it avoids (2)(-2)-class. Our starting point is to develop a similar gluing technique in the setting of branched maximal sections when it does not avoid (2)(-2)-classes. The easiest situation is when the image of the derivatives intersects transversally with the set of planes that are orthogonal to a (2)(-2)-class. More precisely, let $Gr_{3}^{+}(\mathbb{R}^{3,19})$ be the Grassmannian for the positive $3$-plane, and $S_{\delta}=\{H\in Gr^{+}_{3}\big|H\perp\delta\}$. Denote
@@ 2.1. Analysis on flat model
with $|Err(t,z)|\leq Cr^{\frac{5}{2}}\|\rho\|_{,\alpha}$.
@@ 3.1. Non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms on $\mathbb{R}^{n}$
Moreover, the map $(h_{1},\cdots,h_{n-1})\mapsto(a_{1},\cdots,a_{n-1})$ from $(\mathbb{R}_{+})^{n-1}$ to $(\mathbb{R}_{+})^{n-1}$ is surjective.
For completeness, we sketch the construction using Lawlor’s necks. We will begin with a more general scheme of constructing $\mathbb{Z}_{2}$-harmonic $1$-forms from a certain family of special Lagrangians in $\mathbb{C}^{n}$. Regard $T^{*}\mathbb{R}^{n}=(x_{k},y_{k}\frac{\partial}{\partial x_{k}})$ as $\mathbb{C}^{n}$ such that $z_{k}=x_{k}+\sqrt{-1}y_{k}$ is the complex coordinate, and let $(\mathbb{C}^{n},g,\omega,\Omega)$ be the flat Calabi-Yau structure, where $g$ is the Euclidean metric on $\mathbb{C}^{n}$ and
and define a pair of special Lagrangian planes
If there exists a family of special Lagrangians $S_{t}$ such that
  1. (1)
  2. (2)
  3. (3)
  4. (4)
    $S_{t}$ is asymptotic to the $\cup\Pi_{\boldsymbol{\phi}}^{\pm}$ at infinity,
Here, let $C(y^{2})=\prod_{j=1}^{n-1}(1+h_{j}^{-2}y^{2})$, then
Although the above examples are the only known non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms with linear growth and a compact branching set on $\mathbb{R}^{n}$, they are sufficient for the geometric applications in $3$-manifolds, which are our primary focus. In a slightly different direction, based on the uniqueness results for exact special Lagrangians with index $1$ and n1n-1 [17], our examples appear to encompass all exact non-degenerate $\mathbb{Z}_{2}$-harmonic 1-forms with index $1$ and n1n-1.
@@ 3.3. Schauder estimates and elliptic theory
Recall that $\tilde{\alpha}_{\epsilon}$ is the approximate solution constructed in 3.2, and let $\tilde{g}\in\mathcal{M}_{0}^{\epsilon}$ be a small perturbation of the metric. It is easy to check that the error $W^{-1}(\tilde{g})\mathrm{d}^{*}_{\tilde{g}}\tilde{\alpha}_{\epsilon}$ lies in $L^{2}(M\setminus\Sigma_{\epsilon},L_{\epsilon})$. As a result, there is a unique solution to Poisson’s equation
@@ 5.1. Non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-form on compact manifolds with $b^{1}>0$
Remark 5.3.
[Birth of branching sets] Applying the same technique in Proposition 5.1, we can also show that by introducing a large perturbation of the Riemannian metric in a ball, in which $|\alpha|>\delta^{\prime}>0$ for the $\mathbb{Z}_{2}$-harmonic $1$-form $\alpha$, we can thereby construct some new non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms with extra small branching sets contained within the ball.
To prove Proposition 5.1, we will use the following lemma due to Calabi [2]. The idea is to adapt Milnor’s birth lemma of nondegenerate critical points to the context of Calabi’s intrinsic characterization of harmonic $1$-forms on Riemannian manifolds.
@@
[c.f. Section IV, [2]] Let $(M^{n},g)$ be a Riemannian manifold, and let $U$ be an open set in $M^{n}$ on which $\mathrm{d}x$ is a harmonic $1$-form on $U$. There exists a small ball $U^{\prime}\subset U$ such that, in $U^{\prime}$, we can perturb the metric and $1$-form into $(g^{\prime},\mathrm{d}h)$, such that $\mathrm{d}h$ is harmonic with respect to $g^{\prime}$ and has only two regular zeros of index $r,r+1$ for 0<r<n10<r<n-1.
@@
then $f(x_{1}^{\prime},|\boldsymbol{y^{\prime}}|,|\boldsymbol{z^{\prime}}|)$ defines a function with exactly $2$ critical points
@@
Then $f$ is harmonic with respect to the standard Euclidean metric $\sum\mathrm{d}w_{i}^{2}$ on $V_{\pm}$. We can find n2n-2-forms $\xi_{\pm}$ on $B(p_{\pm},3\delta)$ such that
$\mathrm{d}\xi_{\pm}=\star\mathrm{d}f,|\xi_{\pm}|=O(|w|^{2}).$
@@
Let $p\in M^{n}$ and $U$ be sufficiently small neighborhoods of $p$, on which the metric is close to the Euclidean metric. Without loss of generality, we view $U$ as a small ball of radius $3\delta^{\prime}$ in $\mathbb{R}^{n}=(x_{1},\cdots,x_{n})$ and let $\alpha=\mathrm{d}x_{1}$. We similarly define
$\boldsymbol{y}=(x_{2},\cdots,x_{n-r}),\boldsymbol{z}=(x_{n-r+1},\cdots,x_{n}),$
then let $L:\mathbb{R}^{n}\to\mathbb{R}^{n},(x_{1},\cdots,x_{n})\mapsto(Mx_{1}/\delta^{\prime},\cdots,Mx_{n}/\delta^{\prime})$. We perturb $\alpha$ in a ball of radius $\delta^{\prime}$ into
$\beta=\mathrm{d}\bigg(\delta^{\prime}M^{-1}f\circ L\bigg),$
and define $\eta_{\pm}=(\delta^{\prime}/M)^{n-2}L^{*}(\eta_{\pm}^{\prime})$. We also choose a form $\xi_{0}$ such that
@@
and $|\xi_{0}|=O(|x|)$. We define a n2n-2 form
$\tilde{\eta}=\gamma_{1}(|x|/\delta^{\prime})\xi_{0}+\eta_{+}+\eta_{-}.$
By construction, $\beta\wedge\mathrm{d}\xi_{0}>0$ on $\{|x|>2\delta^{\prime}\}$ and in a small punctured ball around each non-degenerate critical point of $f$.
@@
one finds a smooth curve $\ell_{p}$, such that
$p\in\ell_{p},\quad\partial\ell_{p}\subset\partial B(0,3\delta^{\prime}),$
along which $\langle\ell_{p}^{\prime},\beta\rangle>0$. Moreover, we can require that $\ell_{p}$ is the integral curve for the vector field $\frac{\partial}{\partial x_{1}}$ in $B(0,3\delta^{\prime})\setminus B(0,(2+\epsilon)\delta^{\prime})$. For each curve $\ell_{p}$, we define $U_{p}$ to be a tubular neighborhood that does not intersect $L^{-1}(B(p_{\pm},\delta))$ and a map
@@
Let $\xi_{p}$ be a non-negative n1n-1-form on $B^{n-1}(0,1)$, such that $\xi_{p}$ is equal to the Euclidean volume form on $B^{n-1}(0,\frac{1}{3})$ and vanishes outside $B^{n-1}(0,\frac{2}{3})$. It is easy to see that $\beta\wedge\pi_{p}^{*}\xi_{p}>0$ is in a slightly smaller tubular neighborhood. We define $\eta_{p}$ as the cutoff of $\pi_{p}^{*}\xi_{p}$ in the $x_{1}$ direction outside $B(0,2\delta^{\prime})$. Suppose $\{U_{p}\}$ is a finite open cover of $\overline{B(0,2\delta^{\prime})}\setminus L^{-1}B(p_{\pm},2\delta)$, we pick a sufficiently large constant $N$ such that
$\beta\wedge\bigg(\tilde{\eta}+N\sum_{p}\eta_{p}\bigg)>0,$
away from the non-degenerate critical points. By linear algebra, we can find a Riemannian metric $\tilde{g}$ on $B(0,3\delta^{\prime})$ such that $\tilde{g}=g$ near $\partial B(0,3\delta^{\prime})$, $\tilde{g}$ is equal to the Euclidean metric in Morse coordinates near two non-degenerate critical points and $\tilde{\eta}+N\sum_{p}\eta_{p}=\star_{\tilde{g}}\beta$. To conclude, $\beta$ is a harmonic $1$-form with respect to $\tilde{g}$ and agrees with $\mathrm{d}x_{1}$ near $\partial B(0,3\delta^{\prime})$. Extend $\beta$ by $\alpha$ and $\tilde{g}$ by $g$ outside $B(0,3\delta^{\prime})$, we obtain the desired harmonic $1$-form and Riemannian metric. ∎

### v2

@@ Abstract.
We develop a gluing theorem for non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms on compact manifolds, in which non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms on $\mathbb{R}^{n}$ are glued to the regular zeros of a non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-form. As a consequence, we show that for every compact oriented manifold $M^{n},n\geq 3$, if the first Betti number $b^{1}(M)>0$, then $M$ admits a family of non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms, which resolves a folklore conjecture. We will also discuss several possible applications to special holonomy, in particular, to the field of $G_{2}$-geometry.
@@ 1. Introduction
$\alpha=\mathrm{d}\bigg(\mathrm{Re}(A(t)\zeta^{\frac{1}{2}}+B(t)\zeta^{\frac{3}{2}})-\frac{1}{2}\mathrm{Re}(\bar{\mu}\zeta)\mathrm{Re}(A(t)\zeta^{\frac{1}{2}})+o(|\zeta|^{\frac{3}{2}})\bigg),$
@@
The key feature of Theorem 1.3 is that the Riemannian metric g remains unchanged. This is a special feature of dimensions at least 3. In contrast, when M is a compact Riemann surface, the space of $\mathbb{Z}_{2}$-harmonic 1-forms is naturally identified with the finite-dimensional space of quadratic differentials $H^{0}(M,K_{M}^{\otimes 2})$. It follows that, once the metric and therefore the complex structure of the Riemann surface is fixed, there can be no sequence of $\mathbb{Z}_{2}$-harmonic 1-forms whose branching sets is shrinking, equivalently, whose zeros as quadratic differentials collide.
This feature is crucial for geometric applications in special holonomy. On the other hand, if this condition is dropped, one can apply a modification of Calabi’s intrinsic method [2] to construct a family of metrics $g^{\prime}_{\epsilon}$ such that the closed $1$-form obtained in the pre-gluing in Section 3.2 is harmonic with respect to those new metrics. And moreover, in this case, the non-degenerate condition for the $\mathbb{Z}_{2}$-harmonic $1$-form on $\mathbb{R}^{n}$ can be dropped.
@@
The author’s main motivation for studying this problem comes from a non-linear version developed by Donaldson in [9], involving the branched maximal surface equation and the adiabatic limit of Kovalev-Lefchetz coassociative fibration.
Let $\mathbb{R}^{3,19}=H^{2}(K3,\mathbb{R})$ and let $\Lambda=H^{2}(K3,\mathbb{Z})$ be the integral lattice. Also let $\Sigma$ be a smooth link in a compact $3$-manifold $B$ and let $\Gamma$ be the affine extension
@@
$h=(t,\zeta,w,\mathrm{Re}(B(t)\zeta^{\frac{3}{2}}))+o(|\zeta|^{\frac{3}{2}}),$
near $\Sigma$. Here, $t$ is the coordinate on $\Sigma$ and $\zeta$ is the complex coordinate in the normal direction. We call $h$ avoids (2)(-2)-classes if away from the branching set $\Sigma$ there is no (2)(-2)-class orthogonal to the image of the derivatives of $h$. In [9], Donaldson conjectured that if there is a branched maximal section $h$ that avoids (2)(-2)-classes, then there is a compact $7$-manifold $X$ admitting a differentiable Kovalev-Lefschetz fibration $\pi:X\to B$ (Definition 1, page 16 [9]) with generic fibers diffeomorphic to a $K3$ surface, and there is a one parameter family of torsion-free $G_{2}$-structures $\varphi_{t},t>t_{0}\gg 1$ on $X$ such that the fibers of $\pi$ are coassociative. The singular locus of the fibration coincides with $\Sigma$. Moreover, the normalized hyper-symplectic structures on the $K3$ fibers will converge to a family of hyper-Kähler structures induced by the differential of the branched maximal section $h$ at each $b\in B$.
Constructing a solution to the branched maximal surface equation over a compact base space is challenging, particularly under the additional requirement that it avoids (2)(-2)-class. Our starting point is to develop a similar gluing technique in the setting of branched maximal sections when it does not avoid (2)(-2)-classes. The easiest situation is when the image of the derivatives intersects transversally with the set of planes that are orthogonal to a (2)(-2)-class. More precisely, let $Gr_{3}^{+}(\mathbb{R}^{3,19})$ be the Grassmannian for the positive $3$-plane, and $S_{\delta}=\{H\in Gr^{+}_{3}\big|H\perp\delta\}$. Denote
@@
Notation
In this paper, we use $O(|x|^{k})$ to denote the remaining terms for Taylor’s expansion, and $O^{\prime}(|x|^{k})$ to denote the remaining terms such that $|\nabla^{p}f|\leq C_{p}|x|^{k-p}$.
@@ 2.1. Analysis on flat model
with $|Err(t,z)|=o(r^{\frac{3}{2}}\|\rho\|_{,\alpha})$.
@@ 3.1. Non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms on $\mathbb{R}^{n}$
Moreover, the map $(h_{1},\cdots,h_{n-1})\mapsto(a_{1},\cdots,a_{n-1})$ from $(\mathbb{R}_{+})^{n-1}$ to $(\mathbb{R}_{+})^{n-1}$ is bijective.
Although the above examples are the only known non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-forms with linear growth and a compact branching set on $\mathbb{R}^{n}$, they are sufficient for the geometric applications in $3$-manifolds, which are our primary focus. In a slightly different direction, viewing $\mathbb{Z}_{2}$-harmonic $1$-forms as infinitesimal branched deformations of special Lagrangians over $\mathbb{R}^{n}$, based on the uniqueness results for exact special Lagrangians with index $1$ and n1n-1 [17], our examples appear to encompass all exact non-degenerate $\mathbb{Z}_{2}$-harmonic 1-forms with index $1$ and n1n-1.
@@ 3.3. Schauder estimates and elliptic theory
Recall that $\tilde{\alpha}_{\epsilon}$ is the approximate solution constructed in 3.1, and let $\tilde{g}\in\mathcal{M}_{0}^{\epsilon}$ be a small perturbation of the metric. It is easy to check that the error $W^{-1}(\tilde{g})\mathrm{d}^{*}_{\tilde{g}}\tilde{\alpha}_{\epsilon}$ lies in $L^{2}(M\setminus\Sigma_{\epsilon},L_{\epsilon})$. As a result, there is a unique solution to Poisson’s equation
@@ 5.1. Non-degenerate $\mathbb{Z}_{2}$-harmonic $1$-form on compact manifolds with $b^{1}>0$
To prove Proposition 5.1, we will use the following proposition due to Calabi [2]. The idea is to adapt Milnor’s birth lemma of nondegenerate critical points to the context of Calabi’s intrinsic characterization of harmonic $1$-forms on Riemannian manifolds.
@@
[c.f. Section IV, [2]] Let $(M^{n},g)$ be a Riemannian manifold, and let $U$ be an open set in $M^{n}$ on which $\mathrm{d}x$ is a harmonic $1$-form on $U$. There exists a small ball $U^{\prime}\subset U$ such that, in $U^{\prime}$, we can perturb the $1$-form into $\mathrm{d}h$, such that $\mathrm{d}h$ is harmonic with respect to a new Riemannian metric $g^{\prime}$ and has only two regular zeros of index $k,k+1$ for 0<k<n10<k<n-1.
@@
then $\tilde{f}(x_{1}^{\prime},\boldsymbol{y}^{\prime},\boldsymbol{z}^{\prime}):=f(x_{1}^{\prime},|\boldsymbol{y^{\prime}}|,|\boldsymbol{z^{\prime}}|)$ defines a function with exactly $2$ critical points
@@
Then $\tilde{f}$ is harmonic with respect to the standard Euclidean metric $\sum\mathrm{d}w_{i}^{2}$ on $V_{\pm}$. We can find n2n-2-forms $\xi_{\pm}$ on $B(p_{\pm},3\delta)$ such that
$\mathrm{d}\xi_{\pm}=\star\mathrm{d}\tilde{f},|\xi_{\pm}|=O(|w|^{2}).$
@@
Let $p\in M^{n}$ and $U$ be sufficiently small neighborhoods of $p$, on which the metric is close to the Euclidean metric. Without loss of generality, we view $U$ as a small ball of radius $3\delta^{\prime}$ in $\mathbb{R}^{n}=(x_{1},\cdots,x_{n})$ and let $\alpha=\mathrm{d}x_{1}$. Then let $L:\mathbb{R}^{n}\to\mathbb{R}^{n},(x_{1},\cdots,x_{n})\mapsto(Mx_{1}/\delta^{\prime},\cdots,Mx_{n}/\delta^{\prime})$. We perturb $\alpha$ in a ball of radius $\delta^{\prime}$ into
$\beta=\mathrm{d}\bigg(\delta^{\prime}M^{-1}\tilde{f}\circ L\bigg),$
and define $\eta_{\pm}=(\delta^{\prime}/M)^{n-2}L^{*}(\eta_{\pm}^{\prime})$. We also choose a form $\xi_{0}$ in the small ball such that
@@
and $|\xi_{0}|=O(|x|)$. We define a closed n1n-1 form
$\tilde{\zeta}=\begin{cases}\mathrm{d}\bigg(\gamma_{1}(|x|/\delta^{\prime})\xi_{0}+\eta_{+}+\eta_{-}\bigg)&|x|\leq 2\delta^{\prime}\\ \star_{g}\alpha&|x|>2\delta^{\prime}\end{cases}$
By construction, $\beta\wedge\tilde{\zeta}$ on $\{|x|>2\delta^{\prime}\}$ and in a small punctured ball around each non-degenerate critical point of $\beta$.
@@
one finds a smooth curve $\tilde{\ell}_{p}$, such that
$p\in\tilde{\ell}_{p},\quad\partial\tilde{\ell}_{p}\subset\partial B(0,3\delta^{\prime}),$
along which $\langle\tilde{\ell}_{p}^{\prime},\beta\rangle>0$. We can require that $\tilde{\ell}_{p}$ is the integral curve for the vector field $\frac{\partial}{\partial x_{1}}$ in $B(0,3\delta^{\prime})\setminus B(0,(2+\epsilon)\delta^{\prime})$. Moreover, since $\alpha$ is a harmonic $1$-form with respect to the metric $g$, transitivity implies that every two points on $\partial B(0,3\delta^{\prime})$ we can find a smooth curve $\gamma$ on $M\setminus B(0,3\delta^{\prime})$, such that $\langle\gamma^{\prime},\alpha\rangle>0$. Therefore the curve $\tilde{\ell}_{p}$ can be extended smoothly to $\ell_{p}\simeq S^{1}$, with $\langle\ell_{p}^{\prime},\beta\rangle>0$.
For each curve $\ell_{p}$, we define $U_{p}$ to be a tubular neighborhood that does not intersect $L^{-1}(B(p_{\pm},\delta))$ and a map
@@
Let $\xi_{p}$ be a non-negative n1n-1-form on $B^{n-1}(0,1)$, such that $\xi_{p}$ is equal to the Euclidean volume form on $B^{n-1}(0,\frac{1}{3})$ and vanishes outside $B^{n-1}(0,\frac{2}{3})$. It is easy to see that $\beta\wedge\pi_{p}^{*}\xi_{p}>0$ is in a slightly smaller tubular neighborhood. Suppose $\{U_{p}\}$ is a finite open cover of $\overline{B(0,2\delta^{\prime})}\setminus L^{-1}B(p_{\pm},2\delta)$, we pick a sufficiently large constant $N$ such that
$\beta\wedge\bigg(\tilde{\zeta}+N\sum_{p}\pi_{p}^{*}\xi_{p}\bigg)>0,$
away from the non-degenerate critical points. By linear algebra, we can find a Riemannian metric $g^{\prime}$ on $M$ such that $g^{\prime}$ is equal to the Euclidean metric in Morse coordinates near two non-degenerate critical points and $\tilde{\zeta}+N\sum_{p}\pi_{p}^{*}\xi_{p}=\star_{g^{\prime}}\beta$. To conclude, $\beta$ is a harmonic $1$-form with respect to $g^{\prime}$ and agrees with $\alpha$ outside of $B(0,2\delta^{\prime})$ and we obtain the desired harmonic $1$-form and Riemannian metric. ∎